OpenStudy (anonymous):

why the product of any four consecutive integers is divisible by 24

OpenStudy (anonymous):

good question

Directrix (directrix):

From our friends at Wiki: "The product of any four consecutive numbers is divisible by 24. This is because among any four consecutive numbers there must be two even numbers, one of which is a multiple of four, and there must be a multiple of three." Note: Wikipedia is not considered to be an academic research site. http://en.wikipedia.org/wiki/24_(number)

OpenStudy (across):

Let \(a_1\), \(a_2\), \(a_3\), and \(a_4\) be any four consecutive, positive integers. Then we have \(a_1=a_1\), \(a_2=a_1+1\), \(a_3=a_1+2\), and \(a_4=a_1+3\). Multiplying them, we obtain\[a_1a_2a_3a_4=a_1(a_1+1)(a_1+2)(a_1+3)=\]\[a_1^4+6a_1^3+11a_1^2+6a_1.\]From here, you can begin a proof by induction: Notice that if we take \(a_1=1\), the smallest positive integer, we will have \(a_1a_2a_3a_4=24\), which is trivially divisible by \(24\). Can you finish the proof? Or were you just curious about this fact? :)